OpenStudy (anonymous):

a-b=0.9 ab = 0.36 find a+b and a²-b²

ganeshie8 (ganeshie8):

use below : \(\large (a+b)^2 = (a-b)^2 + 4ab\) \(\large a^2-b^2 = (a+b)(a-b)\)

OpenStudy (anonymous):

so is a² + b² = 1.53?????

ganeshie8 (ganeshie8):

why do we need \(a^2 + b^2\) ?

OpenStudy (anonymous):

idk weak at this

ganeshie8 (ganeshie8):

we can find the value of \((a+b)^2\) and take square root to get the value of \(a+b\)

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

a-b=0.9 ab = 0.36 \(\large (a+b)^2 = (a-b)^2 + 4ab = (0.9)^2 + 4(0.36) = ?\)

OpenStudy (anonymous):

its not 4ab its 2ab

OpenStudy (anonymous):

in my country

ganeshie8 (ganeshie8):

haha thats hilarious, whats your country ? :)

OpenStudy (anonymous):

like (a-b)² = a² + b² - 2ab

ganeshie8 (ganeshie8):

yes, add 4ab to that identity and see what you get ^

OpenStudy (anonymous):

?? do and show using 2ab and not 4ab

ganeshie8 (ganeshie8):

(a-b)² + 4ab = a² + b² - 2ab + 4ab = ?

OpenStudy (anonymous):

pl this 4ab is confusing me

OpenStudy (anonymous):

usein the one i yold

OpenStudy (anonymous):

use tthe one i told

ganeshie8 (ganeshie8):

Alright, are these your formulas : \((a+b)^2 = a^2 + b^2 + 2ab\) \((a-b)^2 = a^2 + b^2 - 2ab\) ?

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

subtract them and tell me what you get

OpenStudy (anonymous):

(0.9)² = a² + b² - 2 x 0/36

ganeshie8 (ganeshie8):

that looks good ^

OpenStudy (anonymous):

its .36 not 36

OpenStudy (anonymous):

so now

ganeshie8 (ganeshie8):

find the value of \(a^2 + b^2\)

OpenStudy (anonymous):

1.53????

ganeshie8 (ganeshie8):

Correct !

ganeshie8 (ganeshie8):

\((a+b)^2 = a^2 + b^2 + 2ab\)

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

plugin the values, then take squareroot

OpenStudy (anonymous):

so now what i do

OpenStudy (anonymous):

i alwyas get stuck here

ganeshie8 (ganeshie8):

\((a+b)^2 = 1.53 + 2(0.36)\)

OpenStudy (anonymous):

.81??????

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

tried 8 times

OpenStudy (anonymous):

...

ganeshie8 (ganeshie8):

\((a+b)^2 = 1.53 + 2(0.36) = 2.25\) \(\implies a+b = \pm 1.5\)

OpenStudy (anonymous):

ok now

OpenStudy (anonymous):

a²-b²

ganeshie8 (ganeshie8):

Notice that you got two values for \(a+b\) : -1.5, +1.5

ganeshie8 (ganeshie8):

for calculating \(a^2-b^2\), use the identity : \(a^2 - b^2 = (a+b)(a-b)\)

ganeshie8 (ganeshie8):

when a+b = -1.5 : \(a^2 - b^2 = (-1.5)(0.9) = ?\)

OpenStudy (anonymous):

idk

ganeshie8 (ganeshie8):

just multiply -1.5 and 0.9

ganeshie8 (ganeshie8):

use ur calculator

OpenStudy (anonymous):

more questions

ganeshie8 (ganeshie8):

yea sure

OpenStudy (anonymous):

a-b = 4 a+b = 6 find a² + b² n ab

OpenStudy (anonymous):

is ab = 5?

ganeshie8 (ganeshie8):

you're given system of equations, solve \(a\) and \(b\)

OpenStudy (anonymous):

is ab = 5?

ganeshie8 (ganeshie8):

Yep !

OpenStudy (anonymous):

n the last part 26?

ganeshie8 (ganeshie8):

Correct ! good job :)

OpenStudy (anonymous):

more questions

ganeshie8 (ganeshie8):

okay

OpenStudy (anonymous):

3x + 4y = 16 xy = 4 find the value of 9x² + 16x²

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@skullpatrol

ganeshie8 (ganeshie8):

3x + 4y = 16 square both sides

ganeshie8 (ganeshie8):

(3x + 4y)^2 = 16^2

ganeshie8 (ganeshie8):

expand

ganeshie8 (ganeshie8):

(3x + 4y)^2 = 16^2 (3x)^2 + (4y)^2 + 2(3x)(4y) = 256 9x^2 + 16y^2 + 24xy = 256

ganeshie8 (ganeshie8):

plugin the value of xy

OpenStudy (anonymous):

thtas it?

ganeshie8 (ganeshie8):

yes thats it

OpenStudy (anonymous):

ty

ganeshie8 (ganeshie8):

you're given xy = 4 : 9x^2 + 16y^2 + 24xy = 256 9x^2 + 16y^2 + 24(4) = 256 9x^2 + 16y^2 = 256 - 24(4)

ganeshie8 (ganeshie8):

simplify^